-3x + 5y = 22 (equation 1)
20y − 11 = 12x (equation 2)
This system of linear equations represents________lines.
The system of equation 1 and the equation 20y = 12x + 88 represents______ lines.

Found 2 solutions by Cromlix, MathTherapy:
Answer by Cromlix(4381)

Hi there,
Sorting out into y = mx + c form:-
-3x + 5y = 22
5y = 3x + 22
y = 3/5x + 22/5
.................
20y - 11 = 12x
20y = 12x + 11
y = 12/20x + 11/20
y = 3/5x + 11/20
..................
20y = 12x + 88
y = 12/20x + 88/20
y = 3/5x + 11/5
.................
The system of linear equations (equation 1 and equation 2)
represent parallel lines as the slopes are equal.
Equation 1 and the equation 20y = 12x + 88 represent parallel
lines as both slopes are equal.
Hope this helps :-)

Answer by MathTherapy(10552)


-3x + 5y = 22 (equation 1)
20y − 11 = 12x (equation 2)
This system of linear equations represents________lines.
The system of equation 1 and the equation 20y = 12x + 88 represents______ lines.
The system of equation 1 and the equation 20y = 12x + 88 represents ONE line.